# Badminton Birdie-Like Aerodynamic Alignment of Drifting Dust Grains by Subsonic Gaseous Flows in Protoplanetary Disks

### v1

@@ 1 Introduction
Dust grains within gaseous protoplanetary disks are the raw material for forming planets (e.g. Drążkowska et al., 2023). Their studies have been revolutionized by the Atacama Large Millimeter/submillimeter Array (ALMA), especially through disk-scale observations of linear continuum polarization (e.g. Kataoka et al., 2016b; Stephens et al., 2017; Lee et al., 2018; Girart et al., 2018; Ohashi et al., 2018; Alves et al., 2018; Sadavoy et al., 2018; Bacciotti et al., 2018; Dent et al., 2019; Takahashi et al., 2019; Harrison et al., 2019; Sadavoy et al., 2019; Stephens et al., 2020; Aso et al., 2021; Tang et al., 2023; Stephens et al., 2023; Lin et al., 2024a; Liu et al., 2024; Harrison et al., 2024). The origin of disk polarization is often attributed to scattering when dust grains have grown large enough to efficiently scatter millimeter/sub-millimeter light. Measuring polarization can constrain the properties of dust grains, like the distribution, grain sizes, porosity, etc (e.g. Kataoka et al., 2015; Yang et al., 2016a; Kataoka et al., 2016a; Yang et al., 2017; Kirchschlager & Bertrang, 2020; Lin et al., 2020b; Yang & Li, 2020; Zhang et al., 2023; Zamponi et al., 2024; Yang et al., 2024). Alternatively, if grains are aligned, then the thermal emission is intrinsically polarized (e.g. van de Hulst, 1957; Yang et al., 2016b; Kirchschlager et al., 2019; Lin et al., 2020a, 2022).
@@
In the disk-scales, however, recent observations have revealed compelling evidence for the presence of grains that appear effectively prolate with their long axes aligned toroidally around the disk. For example, multiwavelength observations of HL Tau have unveiled transitions in the polarization morphology from the Very Large Array (VLA) Q-band ($\lambda=7.1$ mm) to ALMA Bands 3, 4, 5, 6, and 7 ranging from $\lambda=$3.1 mm to 0.87 mm (Stephens et al., 2014, 2017; Lin et al., 2024a). The change in pattern cannot be solely attributed to scattering. In particular, Band 3 ($\lambda=$3.1 mm) shows an azimuthally oriented polarization pattern that was shown to require toroidally aligned, effectively prolate grains based on the azimuthal variation in the polarization (Kataoka et al., 2017; Mori & Kataoka, 2021; Yang et al., 2019). Since the optical depth decreases with increasing wavelength, scattering can become less of a factor at longer wavelengths. Multiwavelength models have shown that scattering of toroidally aligned, prolate grains can produce the observed polarization transition (Lin et al., 2022, 2024a). In addition, recent high angular resolution observations ($5$ au resolution) at ALMA Band 7 ($\lambda=0.87$ mm) revealed azimuthally oriented polarization in the first gap, but signatures of scattering polarization in the rings. Stephens et al. (2023) demonstrated that by incorporating optical depth changes between rings and gaps, scattering of toroidally aligned, prolate grains can also explain the observed polarization substructure.
Signatures of toroidally aligned, prolate grains akin to those observed from HL Tau also exist in disks around other sources. Haro 6-13, V892 Tau, DG Tau, and GG Tau show polarization that is predominantly parallel to the disk minor axis (a signature of scattering) at the shorter wavelength, $\lambda\sim 0.9$ mm, and azimuthally oriented polarization at the longer wavelength, $\lambda\sim 3$ mm (Bacciotti et al., 2018; Harrison et al., 2019; Tang et al., 2023; Ohashi et al., 2023; Harrison et al., 2024). For AS 209, the polarization along the disk major axis has polarization angles parallel to the disk minor axis, but the polarization in the outer regions appears azimuthally oriented (at Band 7 from Mori et al. 2019 and at Band 6 from Harrison et al. 2021). The polarization pattern is reminiscent of HL Tau where the optical depth in the inner region is large while the optical depth in the outer region is low leading to a scattering signature in the inner region and a signature of toroidally aligned, prolate grains in the outer region (Stephens et al., 2017; Lin et al., 2024a).
@@
A potential solution is through a modification to the radiative alignment paradigm using the so-called “wrong” internal alignment (Hoang & Lazarian, 2009). For large grains without internal alignment, the short axes can become perpendicular to $\bm{J}$ (in contrast to the “right” alignment mentioned above where the short axes are parallel to $\bm{J}$). When $\bm{J}$ becomes externally aligned to the expected toroidal $\bm{B}$-field of a protoplanetary disk, it can match the desired toroidally aligned prolate grains if grains contain large amounts of iron clusters (Nguyen Tat et al., 2023). However, alignment with the $\bm{B}$-field appears difficult in disk environments because the Larmor precession timescale can be much longer than the gas damping timescale unless the former is greatly shortened by, e.g., superparamagnetic inclusions (Tazaki et al., 2017; Yang, 2021).
@@ 2.1 Problem Setup
Figure 1: Schematic of the double sphere model. The dust grain is composed of two spheres, labeled “1” and “2,” with radii $\varsigma_{1}$ and $\varsigma_{2}$, respectively. Point $O$ is the center of mass of the grain. The radius vectors from the center of mass are $\bm{r}_{1}$ and $\bm{r}_{2}$, respectively. $\bm{A}$ is the gas velocity relative to the center of mass of the grain, i.e., the aerodynamic flow from the perspective of the grain, and is along $\bm{e}_{3}$. The spin of the grain is along $\bm{e}_{2}$. $\theta$ is the angle of $\bm{r}_{2}$ from $\bm{e}_{3}$ in the counter-clockwise direction in this figure. Since alignment occurs at $\theta=0$ when $m_{1}/m_{2}>\varsigma_{1}^{2}/\varsigma_{2}^{2}$, sphere 1 is colored darker to denote a larger mass than sphere 2 and is drawn smaller than sphere 2.
Consider a grain composed of two spheres, sphere 1 and sphere 2, connected by a massless rigid pole. We use point $O$ to denote the center of mass and let $\bm{A}$ be the systematic velocity of the gas with respect to $O$. Fig. 1 shows a schematic of the double-sphere. The radius vectors from $O$ to the center of each spheres are $\bm{r}_{1}$ and $\bm{r}_{2}$, respectively, with corresponding lengths denoted as $r_{1}$ and $r_{2}$. By definition of the center of mass, $r_{i}$ is related to the mass through
$m_{1}r_{1}=m_{2}r_{2}\text{ ,}$ (1)
where $m_{1}$ and $m_{2}$ are the masses of the spheres. The $i$th sphere is characterized by radius $\varsigma_{i}$ and its radius and mass are related through
@@
where $\rho_{i}$ is the material density for the $i$th sphere, which can be different for the spheres.
We use the Cartesian coordinate system centered on the center of mass $O$ with unit directions $\bm{e}_{1}$, $\bm{e}_{2}$, and $\bm{e}_{3}$. Let $\bm{e}_{3}$ be along the direction of the gas flow $\bm{A}$. We define the angle between $\bm{r}_{2}$ and $\bm{e}_{3}$ as $\theta$. Without loss of generality, let $\bm{e}_{1}$ be in the plane formed by $\bm{r}_{2}$ and $\bm{e}_{3}$. The aerodynamic drag force will be in this plane, producing a torque around the $\bm{e}_{2}$ axis, which will be the axis of the flow-induced grain rotating motion, denoted by $\bm{\omega}=\omega\bm{e}_{2}$.
@@
The total torque on the grain is $\bm{\Gamma}=\bm{\Gamma}_{1}+\bm{\Gamma}_{2}$ from both spheres.
We can easily identify that the torque $\bm{\Gamma}$ is only non-zero along $\bm{e}_{2}$ meaning we only have to consider the component $\Gamma_{2}$ ($\Gamma_{1}=\Gamma_{3}=0$; not to be confused with $\bm{\Gamma}_{2}$ which is the torque from the second sphere). There is only one equation of motion:
@@
$D$ $\equiv\frac{4\pi}{3}\rho_{g}v_{\text{th}}(\varsigma_{1}^{2}r_{1}^{2}+\varsigma_{2}^{2}r_{2}^{2})$
$P$ $\equiv\frac{4\pi}{3}\rho_{g}v_{\text{th}}A(-\varsigma_{1}^{2}r_{1}+\varsigma_{2}^{2}r_{2}).$
@@ 2.2 Alignment Direction
where $\theta^{\prime}$ is a dummy variable and the reference potential energy at $\theta=0$ is $-P$. We first consider $P>0$, in which case $U$ is a minimum at $\theta=0$. The $-\cos\theta$ dependence of $U$ is easy to understand through Fig. 2 which shows a schematic of how the orientation of the grain relates to $U$ described by Eq. 13. The direction of the $\theta$-dependent torque depends on the orientation of the grain and acts against the displacement of $\theta$ from 00 attempting to trap the grain in alignment. At $\theta=0$, the torque vanishes and the grain is kept stable with sphere 2 following the direction of the flow (the lighter and/or larger sphere follows $\bm{A}$). The anti-alignment point, $\theta=\pi$, also does not produce a torque, but is unstable. The $\theta$-dependent torque is thus a restoring torque within the potential well, which we call the “flow-induced restoring torque.”
The existence of the damping torque ($\omega$-dependent torque) diminishes the spin, which eventually decreases the rotational energy until $\theta=0$, i.e., the grain becomes aligned, which corresponds to the minimum energy state. Alternatively, if $P<0$ (for example, sphere 2 becomes the heavier sphere), then $U$ reaches a minimum at $\theta=\pi$ though the lighter and/or larger sphere still follows the direction of the flow (recall that $\theta$ is specifically defined using $\bm{r}_{2}$; see Fig. 1).
Figure 2: A schematic of how the orientation of the grain relates to the potential well created by the existence of aerodynamic flow. Top panel: the orientation of an example double-sphere grain model that satisfies $m_{1}/m_{2}>\varsigma_{1}^{2}/\varsigma_{2}^{2}$ like in Fig. 1 The heavier sphere is colored gray, while the lighter sphere is white. The green lines with arrows denote the direction of the gas flow. The circular arrow depicts the direction of the flow-induced torque, which always tries to align the grain and resists any angular displacement. As such, the flow-induced torque is a “restoring" torque. When aligned, the lighter sphere follows the direction of the flow, analogous to how the heavier head of the badminton birdie leads the lighter tail against a headwind. Bottom panel: the energy potential $U$ as a function of $\theta$. The depth of the potential well is $2P$. Alignment occurs at $\theta=0$ where $U$ is minimal.
@@
Observing the expression of $P$ from Eq. 9, we can realize that the quantity $-\varsigma_{1}^{2}r_{1}+\varsigma_{2}^{2}r_{2}$ is the first moment of the cross-sections of the spheres. We can define the “geometrical center” of the double-sphere grain by
$r_{g}\equiv\frac{-\varsigma_{1}^{2}r_{1}+\varsigma_{2}^{2}r_{2}}{\varsigma_{1}^{2}+\varsigma_{2}^{2}}$ (14)
and use $\bm{r}_{g}$ to denote the vector from the center of mass to the geometrical center. When $r_{g}>0$, the geometrical center shifts towards sphere 2, and the final alignment angle at $\theta=0$ is such that $\bm{r}_{g}\parallel\bm{A}$. Conversely, when $r_{g}<0$, the geometrical center moves closer to sphere 1, and the final alignment angle at $\theta=\pi$ also means $\bm{r}_{g}\parallel\bm{A}$. In other words, the asymmetry between the two spheres causes an offset between the center of mass and the geometrical center and the grain becomes susceptible to alignment with $\bm{A}$. The final alignment orientation is such that the geometrical center points along the direction of the flow relative to the center of mass. When $r_{g}=0$, the geometrical center corresponds to the center of mass leading to $P=0$.
@@
$\frac{m_{1}}{m_{2}}>\frac{\varsigma_{1}^{2}}{\varsigma_{2}^{2}}$ (15)
where we used the definition of $P$ and Eq. 1. If the spheres have equal mass, $m_{1}=m_{2}$, then the first sphere should be smaller than the second sphere ($\varsigma_{2}>\varsigma_{1}$) to reach alignment at $\theta=0$.
@@
$\frac{\rho_{1}}{\rho_{2}}\frac{\varsigma_{1}}{\varsigma_{2}}>1.$ (16)
This expression explains that both an asymmetry in the material density and/or in the size can create alignment. If $\rho_{1}=\rho_{2}$, the first sphere needs to be larger than the second sphere $\varsigma_{1}>\varsigma_{2}$ to reach alignment at $\theta=0$ (note the opposite relation because the material density is fixed as opposed to keeping the total mass fixed).

### v2

@@ 1 Introduction
Dust grains within gaseous protoplanetary disks are the raw material for forming planets (e.g. Drążkowska et al., 2023). Their studies have been revolutionized by the Atacama Large Millimeter/submillimeter Array (ALMA), especially through disk-scale observations of linear continuum polarization (e.g. Kataoka et al., 2016b; Stephens et al., 2017; Lee et al., 2018; Girart et al., 2018; Ohashi et al., 2018; Alves et al., 2018; Sadavoy et al., 2018; Bacciotti et al., 2018; Dent et al., 2019; Takahashi et al., 2019; Harrison et al., 2019; Sadavoy et al., 2019; Stephens et al., 2020; Aso et al., 2021; Tang et al., 2023; Stephens et al., 2023; Lin et al., 2024; Liu et al., 2024; Harrison et al., 2024). The origin of disk polarization is often attributed to scattering when dust grains have grown large enough to efficiently scatter millimeter/sub-millimeter light. Measuring polarization can constrain the properties of dust grains, like the distribution, grain sizes, porosity, etc (e.g. Kataoka et al., 2015; Yang et al., 2016a; Kataoka et al., 2016a; Yang et al., 2017; Kirchschlager & Bertrang, 2020; Lin et al., 2020b; Yang & Li, 2020; Zhang et al., 2023; Zamponi et al., 2024; Yang et al., 2024). Alternatively, if grains are aligned, then the thermal emission is intrinsically polarized (e.g. van de Hulst, 1957; Yang et al., 2016b; Kirchschlager et al., 2019; Lin et al., 2020a, 2022).
@@
In the disk-scales, however, recent observations have revealed compelling evidence for the presence of grains that appear effectively prolate with their long axes aligned toroidally around the disk. For example, multiwavelength observations of HL Tau have unveiled transitions in the polarization morphology from the Very Large Array (VLA) Q-band ($\lambda=7.1$ mm) to ALMA Bands 3, 4, 5, 6, and 7 ranging from $\lambda=$3.1 mm to 0.87 mm (Stephens et al., 2014, 2017; Lin et al., 2024). The change in pattern cannot be solely attributed to scattering. In particular, Band 3 ($\lambda=$3.1 mm) shows an azimuthally oriented polarization pattern that was shown to require toroidally aligned, effectively prolate grains based on the azimuthal variation in the polarization (Kataoka et al., 2017; Mori & Kataoka, 2021; Yang et al., 2019). Since the optical depth decreases with increasing wavelength, scattering can become less of a factor at longer wavelengths. Multiwavelength models have shown that scattering of toroidally aligned, prolate grains can produce the observed polarization transition (Lin et al., 2022, 2024). In addition, recent high angular resolution observations ($5$ au resolution) at ALMA Band 7 ($\lambda=0.87$ mm) revealed azimuthally oriented polarization in the first gap, but signatures of scattering polarization in the rings. Stephens et al. (2023) demonstrated that by incorporating optical depth changes between rings and gaps, scattering of toroidally aligned, prolate grains can also explain the observed polarization substructure.
Signatures of toroidally aligned, prolate grains akin to those observed from HL Tau also exist in disks around other sources. Haro 6-13, V892 Tau, DG Tau, and GG Tau show polarization that is predominantly parallel to the disk minor axis (a signature of scattering) at the shorter wavelength, $\lambda\sim 0.9$ mm, and azimuthally oriented polarization at the longer wavelength, $\lambda\sim 3$ mm (Bacciotti et al., 2018; Harrison et al., 2019; Tang et al., 2023; Ohashi et al., 2023; Harrison et al., 2024). For AS 209, the polarization along the disk major axis has polarization angles parallel to the disk minor axis, but the polarization in the outer regions appears azimuthally oriented (at Band 7 from Mori et al. 2019 and at Band 6 from Harrison et al. 2021). The polarization pattern is reminiscent of HL Tau where the optical depth in the inner region is large while the optical depth in the outer region is low leading to a scattering signature in the inner region and a signature of toroidally aligned, prolate grains in the outer region (Stephens et al., 2017; Lin et al., 2024).
@@
A potential solution is through a modification to the radiative alignment paradigm using the so-called “wrong” internal alignment (Hoang & Lazarian, 2009). For large grains without internal alignment, the short axes can become perpendicular to $\bm{J}$ (in contrast to the “right” alignment mentioned above where the short axes are parallel to $\bm{J}$). When $\bm{J}$ becomes externally aligned to the expected toroidal $\bm{B}$-field of a protoplanetary disk, it can match the desired toroidally aligned prolate grains if grains contain large amounts of iron clusters (Thang et al., 2024). However, alignment with the $\bm{B}$-field appears difficult in disk environments because the Larmor precession timescale can be much longer than the gas damping timescale unless the former is greatly shortened by, e.g., superparamagnetic inclusions (Tazaki et al., 2017; Yang, 2021).
@@ 2.1 Problem Setup
Figure 1: Schematic of the double sphere model. The dust grain is composed of two spheres, labeled “black” (circle with a dark shade) and “white” (circle without a shade), with radii $\varsigma_{\bullet}$ and $\varsigma_{\circ}$, respectively. Point $O$ is the center of mass of the grain. The radius vectors from the center of mass to the center of each sphere are $\bm{r}_{\bullet}$ and $\bm{r}_{\circ}$, respectively. The distance between the centers of the spheres is $l$ denoted by a dashed line. $\bm{A}$ is the gas velocity relative to the center of mass of the grain, i.e., the aerodynamic flow from the perspective of the grain, and is along $\bm{\hat{e}}_{3}$. The spin of the grain is along $\bm{\hat{e}}_{2}$. $\theta$ is the angle of $\bm{r}_{\circ}$ from $\bm{\hat{e}}_{3}$ in the counter-clockwise direction in this figure. Since alignment occurs at $\theta=0$ when $m_{\bullet}/m_{\circ}>\varsigma_{\bullet}^{2}/\varsigma_{\circ}^{2}$, the black sphere has a larger mass and is smaller than the white sphere.
Consider a grain composed of two spheres connected by a massless rigid pole. We use point $O$ to denote the center of mass and let $\bm{A}$ be the systematic velocity of the gas with respect to $O$. Fig. 1 shows a schematic of the double-sphere. To distinguish each sphere, we use subscripts $\bullet$” and $\circ$” to represent the black and white spheres, respectively. The radius vectors from $O$ to the center of each sphere are $\bm{r}_{\bullet}$ and $\bm{r}_{\circ}$, respectively, with corresponding lengths denoted as $r_{\bullet}$ and $r_{\circ}$. By definition of the center of mass, the lengths are related to the mass through
$m_{\bullet}r_{\bullet}=m_{\circ}r_{\circ}\text{ ,}$ (1)
where $m_{\bullet}$ and $m_{\circ}$ are the masses of the spheres. The $i$th sphere ($i=\bullet$ or $\circ$) is characterized by radius $\varsigma_{i}$ and its radius and mass are related through
@@
where $\rho_{i}$ is the material density for the $i$th sphere, which can be different for the two spheres.
We use the Cartesian coordinate system centered on the center of mass $O$ with unit directions $\bm{\hat{e}}_{1}$, $\bm{\hat{e}}_{2}$, and $\bm{\hat{e}}_{3}$. Let $\bm{\hat{e}}_{3}$ be along the direction of the gas flow $\bm{A}$. We define the angle between $\bm{r}_{\circ}$ and $\bm{\hat{e}}_{3}$ as $\theta$. Without loss of generality, let $\bm{\hat{e}}_{1}$ be in the plane formed by $\bm{r}_{\circ}$ and $\bm{\hat{e}}_{3}$. The aerodynamic drag force will be in this plane, producing a torque around the $\bm{\hat{e}}_{2}$ axis, which will be the axis of the flow-induced grain rotating motion, denoted by $\bm{\omega}=\omega\bm{\hat{e}}_{2}$.
@@
The total torque on the grain is $\bm{\Gamma}=\bm{\Gamma}_{\bullet}+\bm{\Gamma}_{\circ}$ from both spheres.
We can easily identify that the torque $\bm{\Gamma}$ is only non-zero along $\bm{\hat{e}}_{2}$ meaning we only have to consider the component $\Gamma_{2}$ ($\Gamma_{1}=\Gamma_{3}=0$). There is only one equation of motion:
@@
$D$ $\equiv\frac{4\pi}{3}\rho_{g}v_{\text{th}}(\varsigma_{\bullet}^{2}r_{\bullet}^{2}+\varsigma_{\circ}^{2}r_{\circ}^{2})$
$P$ $\equiv\frac{4\pi}{3}\rho_{g}v_{\text{th}}A(-\varsigma_{\bullet}^{2}r_{\bullet}+\varsigma_{\circ}^{2}r_{\circ}).$
@@ 2.2 Alignment Direction
where $\theta^{\prime}$ is a dummy variable and the reference potential energy at $\theta=0$ is $-P$. We first consider $P>0$, in which case $U$ is a minimum at $\theta=0$. The $-\cos\theta$ dependence of $U$ is easy to understand through Fig. 2 which shows a schematic of how the orientation of the grain relates to $U$ described by Eq. 13. The direction of the $\theta$-dependent torque depends on the orientation of the grain and acts against the displacement of $\theta$ from 00 attempting to trap the grain in alignment. At $\theta=0$, the torque vanishes and the grain is kept stable with the white sphere following the direction of the flow (the less massive and/or larger sphere follows $\bm{A}$). The anti-alignment point, $\theta=\pi$, also does not produce a torque, but is unstable. The $\theta$-dependent torque is thus a restoring torque within the potential well, which we call the “flow-induced restoring torque.”
The existence of the damping torque ($\omega$-dependent torque) diminishes the spin, which eventually decreases the rotational energy until $\theta=0$, i.e., the grain becomes aligned, which corresponds to the minimum energy state. Alternatively, if $P<0$ (for example, the white sphere becomes the heavier sphere), then $U$ reaches a minimum at $\theta=\pi$ though the less massive and/or larger sphere still follows the direction of the flow (recall that $\theta$ is specifically defined using $\bm{r}_{\circ}$; see Fig. 1).
Figure 2: A schematic of how the orientation of the grain relates to the potential well created by the existence of aerodynamic flow. Top panel: the orientation of an example double-sphere grain model that satisfies $m_{\bullet}/m_{\circ}>\varsigma_{\bullet}^{2}/\varsigma_{\circ}^{2}$ like in Fig. 1 The heavier sphere is colored gray, while the less massive sphere is white. The green lines with arrows denote the direction of the gas flow. The circular arrow depicts the direction of the flow-induced torque, which always tries to align the grain and resists any angular displacement. As such, the flow-induced torque is a “restoring" torque. When aligned, the less massive sphere follows the direction of the flow, analogous to how the heavier head of the badminton birdie leads the less massive tail against a headwind. Bottom panel: the energy potential $U$ as a function of $\theta$. The depth of the potential well is $2P$. Alignment occurs at $\theta=0$ where $U$ is minimal.
@@
Observing the expression of $P$ from Eq. 9, we can realize that the quantity $-\varsigma_{\bullet}^{2}r_{\bullet}+\varsigma_{\circ}^{2}r_{\circ}$ is the first moment of the cross-sections of the spheres. We can define the “geometrical center” of the double-sphere grain by
$r_{g}\equiv\frac{-\varsigma_{\bullet}^{2}r_{\bullet}+\varsigma_{\circ}^{2}r_{\circ}}{\varsigma_{\bullet}^{2}+\varsigma_{\circ}^{2}}$ (14)
and use $\bm{r}_{g}$ to denote the vector from the center of mass to the geometrical center. When $r_{g}>0$, the geometrical center shifts towards the white sphere and the final alignment angle at $\theta=0$ is such that $\bm{r}_{g}\parallel\bm{A}$. Conversely, when $r_{g}<0$, the geometrical center moves closer to the black sphere, and the final alignment angle at $\theta=\pi$ also means $\bm{r}_{g}\parallel\bm{A}$. In other words, the asymmetry between the two spheres causes an offset between the center of mass and the geometrical center and the grain becomes susceptible to alignment with $\bm{A}$. The final alignment orientation is such that the geometrical center points along the direction of the flow relative to the center of mass. When $r_{g}=0$, the geometrical center corresponds to the center of mass leading to $P=0$.
@@
$\frac{m_{\bullet}}{m_{\circ}}>\frac{\varsigma_{\bullet}^{2}}{\varsigma_{\circ}^{2}}$ (15)
where we used the definition of $P$ and Eq. 1. If the spheres have equal mass, $m_{\bullet}=m_{\circ}$, then the black sphere should be smaller than the white sphere ($\varsigma_{\circ}>\varsigma_{\bullet}$) to reach alignment at $\theta=0$.
@@
$\frac{\rho_{\bullet}}{\rho_{\circ}}\frac{\varsigma_{\bullet}}{\varsigma_{\circ}}>1.$ (16)
This expression explains that both an asymmetry in the material density and/or in the size can create alignment. If $\rho_{\bullet}=\rho_{\circ}$, the black sphere needs to be larger than the white sphere $\varsigma_{\bullet}>\varsigma_{\circ}$ to reach alignment at $\theta=0$ (note the opposite relation because the material density is fixed as opposed to keeping the total mass fixed).